One pump can fill a swimming pool in 8 hours and another pump can fill it in 10 hours. If both pumps are opened at the same time, how many hours will it take to fill the pool?

Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

Let x be the amount of time it takes to fill the pool with both pumps working together, then
:
(1/8) + (1/10) = (1/x)
:
represents amount of pool filled in an hour working together
:
(5/40) + (4/40) = (1/x)
:
9/40 = 1/x
:
9x = 40
:
x = 4.44444 approx 4.4 hours
:
******************************************
it will take 4.4 hours to fill the pool
******************************************
:

Answer by ikleyn(52781)

For a bunch of similar solved joint work problems see the lessons
    - Using Fractions to solve word problems on joint work,
    - Solving more complicated word problems on joint work,
    - Selected joint-work word problems from the archive
in this site.

Read them and become an expert in solving these problems.